# Do Unlearning Methods Remove Information from Language Model Weights?

### v2

@@ Abstract
Large Language Models’ knowledge of how to perform cyber-security attacks, create bioweapons, and manipulate humans poses risks of misuse. Previous work has proposed methods to unlearn this knowledge. Historically, it has been unclear whether unlearning techniques are removing information from the model weights or just making it harder to access. To disentangle these two objectives, we propose an adversarial evaluation method to test for the removal of information from model weights: we give an attacker access to some facts that were supposed to be removed, and using those, the attacker tries to recover other facts from the same distribution that cannot be guessed from the accessible facts. We show that using fine-tuning on the accessible facts can recover 88% of the pre-unlearning accuracy when applied to current unlearning methods, revealing the limitations of these methods in removing information from the model weights. footnotetext: Code available at https://github.com/aghyad-deeb/unlearning_evaluation
@@ 1 Introduction
To address these concerns, LLMs are typically trained to refuse to engage in dangerous activities. Refusal is vulnerable to jailbreak techniques (Wei et al., 2023; Zou et al., 2023; Liu et al., 2024b) and other attacks. We can address these vulnerabilities by ensuring that dangerous knowledge is not present in the weights. Filtering out dangerous knowledge from the training data of LLMs and rerunning pretraining is impractical given the size of the pretraining datasets. Machine unlearning was suggested to remove harmful knowledge from models (Si et al., 2023; Li et al., 2024b), offering a stronger safety assurance relative to refusal.
@@
    We run evaluations on common unlearning methods. This includes Gradient Ascent, RMU, and training on incorrect facts. We show that after performing our attack to recover hidden information, we can recover at least 88% of the pre-unlearning accuracy for all the unlearning methods we evaluate when the unlearning maintains good performance on non-unlearned tasks.
@@ 2 Related Work
Past work has introduced white-box unlearning evaluations like linear probes and relearning time. Linear probes may recover the information present in the activations (Lynch et al., 2024b), but are not powerful enough to detect information present in the weights. For example, probes on top of RMU models fail to get high accuracy, but RMU models can still be jailbroken. Relearning time is a promising approach to evaluate unlearning that was used by Golatkar et al. (2020a), Golatkar et al. (2020b), Tarun et al. (2023), and Lynch et al. (2024b). This approach measures the number of fine-tuning steps needed to recover accuracy on the unlearned knowledge, where the fine-tuning data is drawn from the dataset we perform unlearning on. If accuracy is recovered after a small number of fine-tuning steps, this is evidence that the unlearning technique is ineffective at removing knowledge. The main problem with this approach is that it produces a metric that is only meaningful when compared to the learning time for a model that was never trained on the knowledge in the first place, which is hard to obtain given the cost of LLM pretraining. This method also suffers against methods that specifically target making fine-tuning slower or more difficult (Henderson et al., 2023; Rosati et al., 2024; Tamirisa et al., 2024).
@@
Previous work has shown evidence about current unlearning techniques being weak against attacks that would fail if the information was removed (Lynch et al., 2024a; Łucki et al., 2024; Hong et al., 2024). Our results further confirm the findings of this previous work and provides a standard method that can be applied to any unlearned model (a model to which unlearning was applied). Further, negative results using our method provide strong evidence of unsuccessful knowledge removal relative to previous prompting-based methods.
@@ 3.2 Estimating the Presence of Information
Because unlearning was performed on $T$ and $V$, access to $T$ allows for the creation of attacks that can revert the hiding behavior that some unlearning methods may lead to in the model, and because the facts in $T$ and $V$ are independent, we do not need to worry about “reteaching” the model the facts, which is a concern if we perform attacks that use access to $V$, like relearning time (see Section 2).
The developers of the unlearning method try to find an unlearning technique $U^{*}$ that minimizes the recovered accuracy on $V$:
@@ 4.1 Datasets
We create several datasets that differ in how much they fulfill each of these properties. For each of these datasets, we also have retain datasets that unlearning methods use to ensure the model does not unlearn capabilities we want it to keep:
@@
  • Random Birthdays: A dataset with randomly generated names and randomly generated years of birth. As it is randomly generated, we first fine-tune the models on the dataset, unlike the other 3 datasets. We use 4 splits for T and 1 split for V. We use a subset of the MMLU categories for the retain dataset. The creation of the Random Birthdays dataset was inspired by Maini et al. (2024), and we use it to test unlearning methods when we are confident that the facts are independent. We test that the facts are indeed independent and show the results in Appendix E.
@@ 4.2 Unlearning
    Gradient Difference (GD) (Liu et al., 2022): Gradient Ascent on the forget dataset and Gradient Descent on the retain dataset. $\text{Loss}_{\text{GD}}=-\text{Loss}_{\text{Forget}}+\alpha*\text{Loss}_{\text{Retain}}$ Where $\alpha$ is the retain coefficient.
@@ 5 Results
Figure 2: Forget accuracies before and after RTT for different unlearning methods and Datasets. We perform unlearning using RMU, GD, and RIA then perform RTT. The unlearning strength is chosen such that the drop in the retain accuracy is less than or equal to 5%, where the unlearning strength is controlled by adjusting the corresponding hyperparameter (see Section 4.2) in each unlearning method. The results for a retain accuracy drop of less than or equal to 10%, 30% and 100% are available in Appendix G
@@
As shown in Figure 3, we find that both RMU and GD successfully reduce the accuracy after performing unlearning. RIA leads to less significant reductions in accuracy. For all methods, RTT recovers the forget accuracy close to its original level, which suggests that most of the information was hidden, not removed from the weights.
@@
A lower recovery rate corresponds to more successful information removal. In our tests, all recovery rates were greater than 88%, implying poor performance at removing information.
To test whether the retain loss is restricting unlearning methods from appropriately removing the information from the weights, we run unlearning with different unlearning strengths to achieve different values for the retain accuracy. Figure 3 shows that even with large losses in the retain accuracy, RTT is able to recover accuracy on the forget dataset. Even if we don’t include a retain loss, RTT is often able to recover forget accuracy (see Appendix G, Figure 10). RTT recovering accuracy even when the model is not incentivized to retain performance on other tasks implies that the unlearning methods are not restricted by having to maintain good performance on the retain dataset.
Figure 4: Forget accuracies for different formats of the unlearning dataset. We perform unlearning and RTT for different text formats and loss types when using RMU and GD. The unlearning strength is such that the loss in the retain accuracy is less than or equal to 5%. All of the runs were done using the WMDP-Deduped dataset.
We test how the format of text used for unlearning affects performance. The results are shown in Figure 4. Using the plain-text generated by GPT-4o provides the best balance of performance across different unlearning methods, in addition to being generalizable to all MCQ datasets. RMU performs better than GD when the unlearning dataset is related but doesn’t necessarily contain the same facts as the ones used in RTT and evaluation. GD performs best when all of the unlearning dataset, RTT and evaluation use the MCQ format and the loss is restricted to the answer tokens. These observations may imply that RMU tends to generalize the unlearning more than GD does.
@@ 6.1 High Granularity Knowledge Hiding
We perform unlearning using GD with multiple hyperparameters and loss types. We discuss the hyperparameters in Appendix H. With the hyperparameters that obtain the best trade-off between forget accuracy and retain accuracy, we were able to obtain was 32% forget accuracy and 72% retain accuracy. We then perform RTT on the model to see if we can recover the lost accuracy. We were able to recover the forget accuracy back to 92%.
@@ 7.2 Recommendations
In the light of our work and inspired by Carlini et al. (2019), we make the following recommendations for future research on addressing dangerous capabilities and knowledge of Artificial Intelligence models:
@@
    Indicate whether the purpose of the proposed method is to remove information or make the information harder to access in the model.
@@
    When the goal is removing capabilities and/or knowledge, evaluate the proposed method against attacks that aim to recover them, like using the RTT attack presented in this paper.
@@ Appendix A RTT Hyperparameters
$2\sigma=1.96\frac{\sqrt{p(1-p)}}{n}\leq 1.96\,\sqrt{\frac{1}{4n}}$
@@ Appendix C Fine-tuning Robust Methods
Other techniques include Tampering Attack Resistance (TAR) introduced by Tamirisa et al. (2024), but this technique is vulnerable against parameter-efficient fine-tuning (PEFT) as demenstorated in the work.
Overall, because fine-tuning robustness techniques can be bypassed when using an extensive hyperparameter search, we think using RTT with an extensive hyperparameter search would still expose knowledge that wasn’t removed.
@@ Appendix G Results for Different Drops in Retain Accuracies
Appendix H Stress-Testing Hyperparameters
When performing the stress-test described in section 6, we first fine-tune the original model on the Random Birthdays dataset while freezing the second half of the model. We then perform unlearning using GD while freezing the first half of the model. Performing this unlearning required an extensive hyperparameter search. We end up using the Lion optimizer (Chen et al., 2023), a 6.4e-6 learning rate, 11 epochs, and a retain coefficient of 1. In addition, we restrict the loss to be on the number tokens, to help the model focus on the tokens that contain information that should be unlearned.
@@ Appendix I WMDP vs WMDP-Deduped
We create WMDP-Deduped to create a version of WMDP that has low shared information among facts. We do this by removing skill-based questions (which use the same templates and rely on similar facts) using keyword matching (e.g. ”What is the output”), and removing duplicate questions using the cosine similarity of questions and answers separately, using OpenAI’s text-embedding-3-large. We use an aggressive threshold of 0.7 for WMDP-cyber and 0.6 for WMDP-bio. We manually check that pairs of questions just below these thresholds don’t seem to be about the same facts.
We don’t include any of the 408 questions from WMDP-chem both for simplicity and because the original WMDP paper focuses on unlearning in cybersecurity and biosecurity but not in chemistry.

### v3

@@ Abstract
Large Language Models’ knowledge of how to perform cyber-security attacks, create bioweapons, and manipulate humans poses risks of misuse. Previous work has proposed methods to unlearn this knowledge. Historically, it has been unclear whether unlearning techniques are removing information from the model weights or just making it harder to access. To disentangle these two objectives, we propose an adversarial evaluation method to test for the removal of information from model weights: we give an attacker access to some facts that were supposed to be removed, and using those, the attacker tries to recover other facts from the same distribution that cannot be guessed from the accessible facts. We show that using fine-tuning on the accessible facts can recover 88% of the pre-unlearning accuracy when applied to current unlearning methods for information learned during pretraining, revealing the limitations of these methods in removing information from the model weights. Our results also suggest that unlearning evaluations that measure unlearning robustness on information learned during an additional fine-tuning phase may overestimate robustness compared to evaluations that attempt to unlearn information learned during pretraining. footnotetext: Code available at https://github.com/aghyad-deeb/unlearning_evaluation
@@ 1 Introduction
To address these concerns, LLMs are typically trained to refuse to engage in dangerous activities. Refusal is vulnerable to jailbreak techniques (Wei et al., 2023; Zou et al., 2023; Liu et al., 2024b) and other attacks. We can address these vulnerabilities by ensuring that dangerous knowledge is not present in the weights. Filtering out dangerous knowledge from the training data of LLMs and rerunning pretraining is impractical given the size of the pretraining datasets. Machine unlearning was suggested to remove harmful knowledge from models (Si et al., 2023; Li et al., 2024b; Barez et al., 2025), offering a stronger safety assurance relative to refusal.
@@
    We run evaluations on common unlearning methods. This includes Gradient Ascent, RMU, and training on incorrect facts. We show that after performing our attack to recover hidden information, we can recover at least 88% of the pre-unlearning accuracy for all the unlearning methods we evaluate when unlearning information learned in pretraining, and when the unlearning maintains good performance on non-unlearned tasks.
  3. 3.
    We compare the robustness of unlearning information learned during pretraining (the primary use case of unlearning) against unlearning information learned during an additional fine-tuning phase (a method sometimes used in unlearning evaluations).
  4. 4.
    We stress-test our framework against a harder case for recovering hidden knowledge.
@@ 2 Related Work
Threat model Metric What is being measured Papers
Attacks that do not require knowledge of the unlearned information: jailbreaks, steering vectors, etc. Accuracy on V after a medium-scale retraining on T  [^1] The terms T and V are introduced in Section 3.2. Is the information still present in the weights such that a medium-scale fine-tuning attack can recover it? Ours
Accuracy after a small-scale relearning attack Is the information still present in the weights such that a small-scale fine-tuning attack can recover it? Hu et al. (2024), Łucki et al. (2024)
Success rate of a sample of attacks that do not require knowledge of the unlearned information Are some of these attacks successful? Lynch et al. (2024a), Li et al. (2024b)
Relearning attacks: getting the model to output unlearned information with limited computational resources Relearning time, relearning sample efficiency Is it possible to cheaply make the model useful again at the unlearned task? Tamirisa et al. (2024b)
Regular usage (unintentional attacks) Direct prompting How often is unlearned information unintentionally revealed? Similar to Carlini et al. (2019b)
Table 1: A comparison of the target threat model, the used metric, and what is being measured in different approaches for evaluating unlearning.
@@
Past work has introduced white-box unlearning evaluations like linear probes and relearning time. Some of the white-box appraoches include linear probes and relearning with fine-tuning. Linear Probes may recover the information present in the activations (Lynch et al., 2024b), but are not powerful enough to detect information present in the weights. For example, probes on top of RMU models fail to get high accuracy, but RMU models can still be jailbroken. Relearning time and limited-sample relearning is a promising approach to evaluate unlearning that was used by Golatkar et al. (2020a), Golatkar et al. (2020b), Tarun et al. (2023), and Lynch et al. (2024b). Relearning time measures the number of fine-tuning steps needed to recover accuracy on the unlearned knowledge, and limited-sample relearning fine-tunes on a limited number of samples, where the fine-tuning data is drawn from the dataset we perform unlearning on. If accuracy is recovered after a small number of fine-tuning steps or with a limited number of samples, this is evidence that the unlearning technique is ineffective at removing knowledge. The main problem with this approach is the lack of a reliable baseline for comparison. if the information is recovered after a specific number of fine-tuning steps or with a specific number of training samples, we cannot determine if this number of time steps or samples implies that the information was hidden or removed. This method also suffers against methods that specifically target making fine-tuning slower or more difficult (Henderson et al., 2023; Rosati et al., 2024; Tamirisa et al., 2024a). Table 1 shows previous approaches and how our approach compares. We expect the results from other fine-tuning-based approaches that test for the existence of hidden knowledge to be highly correlated with ours, but expect our approach to have an advantage. Specifically, given the design of our dataset where the facts are independent, we can perform a larger-scale retraining/fine-tuning. This is more challenging in other approaches, as they use facts from the same dataset, which means they either fine-tune on a small number of facts to ensure no leakage happens, or they use more facts, which risks leakage happening.
@@
Previous work has shown evidence about current unlearning techniques being weak against attacks that would fail if the information was removed (Lynch et al., 2024a; Łucki et al., 2024; Hong et al., 2024). Our results further confirm the findings of this previous work. Further, negative results using our method provide strong evidence of unsuccessful knowledge removal relative to previous prompting-based methods.
@@ 3.2 Estimating the Presence of Information
Because unlearning was performed on $T$ and $V$, access to $T$ allows for the creation of attacks that can revert the hiding behavior that some unlearning methods may lead to in the model, and because the facts in $T$ and $V$ are independent, we do not need to worry about “reteaching” the model the facts, which is a concern if we perform attacks that use access to $V$, like relearning time (see Section 2). The developers of the unlearning method try to find an unlearning technique $U^{*}$ that minimizes the recovered accuracy on $V$:
@@ 4.1 Datasets
We distinguish between two types of datasets: the forget dataset and the retain dataset. The forget dataset consists of questions that correspond to facts we want the model to unlearn. The retain dataset consists of questions that correspond to facts that we would not want the model to unlearn; we want the model to still know these facts after unlearning. We create several forget datasets that vary in how well they fulfill each of the properties mentioned above and match them with appropriate retain datasets:
@@
  • Random Birthdays (fine-tuned information): A dataset with randomly generated names and randomly generated years of birth. As it is randomly generated, we first fine-tune the models on the dataset, unlike the other 3 datasets. We use 4 splits for T and 1 split for V. We use a subset of the MMLU categories for the retain dataset. The creation of the Random Birthdays dataset was inspired by Maini et al. (2024), and we use it to test unlearning methods when we are confident that the facts are independent. We test that the facts are indeed independent and show the results in Appendix E.
@@ 4.2 Unlearning
    Gradient Difference (GD) (Liu et al., 2022): Gradient Ascent on the forget dataset and Gradient Descent on the retain dataset. $\text{Loss}_{\text{GD}}=-\text{Loss}_{\text{Forget}}+\alpha*\text{Loss}_{\text{Retain}}$ Where $\alpha$ is the retain coefficient. We note that using a retain coefficient of 00 corresponds to using the Gradient Ascent unlearning method.
@@ 5 Results
Figure 2: Forget accuracies before and after RTT for different unlearning methods and Datasets. We perform unlearning using RMU, GD, and RIA then perform RTT. The unlearning strength is chosen such that the drop in the retain accuracy is less than or equal to 5%, where the unlearning strength is controlled by adjusting the corresponding hyperparameter (see Section 4.2) in each unlearning method. The results for a retain accuracy drop of less than or equal to 10%, 30% and 100% are available in Appendix G.
@@
5.1 Unlearning Pretrained Information
As shown in Figure 2, we find that both RMU and GD successfully reduce the accuracy after performing unlearning. RIA leads to less significant reductions in accuracy. For all methods, RTT recovers the forget accuracy close to its original level for pretrained information (Years, MMLU, and WMDP-Deduped datasets), which suggests that most of the information was hidden, not removed from the weights.
@@ 5.1 Unlearning Pretrained Information
A lower recovery rate corresponds to more successful information removal. In our tests, recovery rates for pretrained information were greater than 88%, implying poor performance at removing information.
To test whether the retain loss is restricting unlearning methods from appropriately removing the information from the weights, we run unlearning with different unlearning strengths to achieve different values for the retain accuracy. Figure 3 shows that even with large losses in the retain accuracy, RTT is able to recover accuracy on the forget dataset with pretrained information. Even if we do not include a retain loss, RTT is often able to recover forget accuracy (see Appendix G, Figure 10). RTT recovering accuracy even when the model is not incentivized to retain performance on other tasks implies that the unlearning methods are not restricted by having to maintain good performance on the retain dataset.
5.2 Unlearning Fine-tuned Information
RTT does not recover information for the fine-tuned dataset (Random Birthdays); we still get a high recovery rate for RMU, but a recovery rate less than $35\%$ for GD and RIA. We suspect that changes introduced by fine-tuning might be relatively superficial and easy to reverse as observed by Jain et al. (2024). However, we only experiment with one dataset with fine-tuned information and leave it to future work to determine how easy it is to robustly remove information learned by fine-tuning.
Figure 4: Forget accuracies for different formats of the unlearning dataset. We perform unlearning and RTT for different text formats and loss types when using RMU and GD. The unlearning strength is such that the loss in the retain accuracy is less than or equal to 5%. All of the runs were done using the WMDP-Deduped dataset. ( For ”MCQ with Loss on Answer Only”, we do not consider RMU as RMU acts on activations of intermediate layers and we cannot restrict the loss on answer tokens.)
5.3 Format of Unlearning Datasets
We test how the format of text used for unlearning affects performance. The results are shown in Figure 4. Using the plain-text generated by GPT-4o (examples can be found in Appendix K) provides the best balance of performance across different unlearning methods, in addition to being generalizable to all MCQ datasets. RMU performs better than GD when the unlearning dataset is related but does not necessarily contain the same facts as the ones used in RTT and evaluation. GD performs best when all of the unlearning dataset, RTT and evaluation use the MCQ format and the loss is restricted to the answer tokens. These observations may imply that RMU tends to generalize the unlearning more than GD does.
@@ 6.1 High Granularity Knowledge Hiding
We perform unlearning using GD with multiple hyperparameters and loss types. We discuss the hyperparameters and additional details in Appendix H. With the hyperparameters that obtain the best trade-off, we were able to obtain 35% forget accuracy and 82% retain accuracy. We then perform RTT on the model to see if we can recover the lost accuracy. We were able to recover the forget accuracy back to 83% quickly.
@@ 7.2 Recommendations
In the light of our work and inspired by Carlini et al. (2019a), we make the following recommendations for future research on addressing dangerous capabilities and knowledge of Artificial Intelligence models:
@@
    Indicate whether the purpose of the proposed method is to remove inforLmation or make the information harder to access in the model.
@@
    When the goal is removing capabilities and/or knowledge, evaluate the proposed method against attacks that aim to recover them, like using the RTT attack presented in this paper. Additionally, we recommend testing against information learned during pretraining, as information learned by fine-tuning might not transfer to realistic settings; our evaluations on the Random Birthdays dataset overestimate the robustness of some unlearning techniques compared to datasets containing pretrained information.
@@ Appendix A RTT Hyperparameters
$2\sigma=1.96\sqrt{\frac{p(1-p)}{n}}\leq 1.96\,\sqrt{\frac{1}{4n}}$
@@ Appendix C Fine-tuning Robust Methods
Other techniques include Tampering Attack Resistance (TAR) introduced by Tamirisa et al. (2024a), but this technique is vulnerable against parameter-efficient fine-tuning (PEFT) as demenstorated in the work.
Overall, because fine-tuning robustness techniques can be bypassed when using an extensive hyperparameter search, we think using RTT with an extensive hyperparameter search would still expose knowledge that was not removed.
@@ Appendix G Results for Different Drops in Retain Accuracies
Appendix H Stress-Testing Details
When performing the stress-test described in section 6, we first fine-tune the original model on the Random Birthdays dataset while freezing the second half of the model. We then perform unlearning using GD while freezing the first half of the model. Performing this unlearning required an extensive hyperparameter search and multiple trials with the final hyperparameter configuration to get unlearning to wor. In runs where unlearning succeeds, we use the Lion optimizer (Chen et al., 2023), a 6.4e-6 learning rate, 11 epochs, and a retain coefficient of 1. In addition, we restrict the loss to be on the number tokens, to help the model focus on the tokens that contain information that should be unlearned. While unlearing took many trials to succeed, RTT recovers information successfully without multiple trials.
@@ Appendix I WMDP vs WMDP-Deduped
We create WMDP-Deduped to create a version of WMDP that has low shared information among facts. We do this by removing skill-based questions (which use the same templates and rely on similar facts) using keyword matching (e.g. ”What is the output”), and removing duplicate questions using the cosine similarity of questions and answers separately, using OpenAI’s text-embedding-3-large. We use an aggressive threshold of 0.7 for WMDP-cyber and 0.6 for WMDP-bio. We manually check that pairs of questions just below these thresholds do not seem to be about the same facts.
We do not include any of the 408 questions from WMDP-chem both for simplicity and because the original WMDP paper focuses on unlearning in cybersecurity and biosecurity but not in chemistry.
@@
Appendix L Historical Note
In an earlier version of this preprint, we used a version of the Random Birthdays that had a few duplicate names; the same name appeared more than once with different birthdays. This was not our intention. When using this old version, we were able to recover the information after unlearning using RTT.
Thanks to an ICLR reviewer who asked about name duplicates, we noticed the issue and fixed it.
The current version of the Random Birthdays dataset has no duplicates. Removing duplicates and replacing them with other unique name-birthday pairs changed the results, and RTT cannot recover most information after unlearning with GD and RIA.